### v3

# A Differentiable Sparse Vector Quantization (SVQ) for Spatio-Temporal Forecasting

Abstract
Spatio-temporal forecasting, pivotal in numerous fields, hinges on the delicate equilibrium between isolating nuanced patterns and sifting out noise. To tackle this, we introduce Sparse Regression-based Vector Quantization (SVQ), a novel technique that leverages sparse regression for succinct representation, an approach theoretically and practically favored over classical clustering-based vector quantization methods. This approach preserves critical details from the original vectors using a regression model while filtering out noise via sparse design. Moreover, we approximate the sparse regression process using a blend of a two-layer MLP and an extensive codebook. This approach not only substantially cuts down on computational costs but also grants SVQ differentiability and training simplicity, resulting in a notable enhancement of performance. Our empirical studies on five spatial-temporal benchmark datasets demonstrate that SVQ achieves state-of-the-art results. Specifically, on the WeatherBench-S temperature dataset, SVQ improves the top baseline by 7.9%. In video prediction benchmarks—Human3.6M, KTH, and KittiCaltech—it reduces MAE by an average of 9.4% and improves image quality by 17.3% (LPIPS). Code is publicly available at https://anonymous.4open.science/r/SVQ-Forecasting.
1 Introduction
Spatio-temporal forecasting (Wang et al., 2022a) has board applications in many domains including environment, climate, transportation, etc. Many traditional methods relying on feature engineering have been studied in the past several decades. Recently, deep learning models are becoming more popular (Wang et al., 2022a). However, as more applications (e.g., weather forecasting) arise, it calls for deep models to accurately predict future dynamics with precise spatial and temporal resolutions.
Despite the various forms of spatio-temporal forecasting approaches, such as the use of graph neural network (GNN) models for unevenly resolved data, our work concentrates on grid-based datasets, one of the most widely adopted structures in the field. Within this context, we address a primary challenge: how to finely balance the preservation of important details against the need to filter out irrelevant noise, thereby enhancing the generalized performance of forecasting models. Although vector quantization (VQ) is effective in reducing noises by constraining latent patterns to a clustered codebook, it is usually limited in preserving the diversity of input patterns, making it suboptimal for spatial-temporal forecasting. We performed a comprehensive suite of experiments, the results of which indicate that, contrary to expectations, nearly all sota VQ methods we examined did not enhance the accuracy of a basic weather forecasting model as shown in Table 4. This outcome may provide insight into the infrequent application of VQ techniques in spatial-temporal prediction tasks.
To address this limitation of VQ, we propose a Sparse regression-based Vector Quantization (SVQ) method. By associating each vector to multiple codes from a large codebook, it is able to cover diverse visual patterns; by introducing a sparse regularization mechanism in the regression model, we effectively filter out noise and thus ensure a good generalized performance. As shown in Figure 1, our proposed SVQ method exhibits a more uniform distribution of quantized vectors compared to VQ, indicating that SVQ is able to preserve more diverse and fine-grained information from the original inputs than VQ.
Despite employing a series of vectors for representation instead of a single codebook vector, regression-based vector quantization remains true to the foundational aspects of the technique. It still quantizes input vectors into a finite set of discrete levels, utilizing a dynamic codebook-like structure. The key process of mapping input data to representative vectors is maintained, but with the added sophistication of regression models that adaptively combine vectors for a more accurate representation. This method preserves the essence of vector quantization—efficiently compressing and approximating data—while enhancing its ability to capture complex patterns within the input space.
To address the high computational cost of sparse regression, which scales quadratically with codebook size, we approximate the sparse regression model with a MLP model and a large codebook. It not only dramatically simplifies the computation, but more importantly, makes SVQ differentiable and thus easy to train. We emphasize the importance of differentiability of SVQ for its success in spatial-temporal forecasting. Our experiments reveal that lookup-based VQ methods can substantially increase prediction error, despite some employing a residual design that theoretically should perform on par with or better than non-VQ cases. This suggests the “straight-through” estimator, a non-differentiable component of VQ, may be a major factor behind VQ’s underperformance. In particular, the non-differentiable operator such as argmin or lookup introduces errors into gradient estimation of all modules preceding the quantization module, consequently undermining prediction accuracy. In contrast, our proposed SVQ is differentiable and easy to compute accurate gradients for training. We finally note that SVQ also differs from traditional VQ methods by not relying on intricate models to construct the prior distribution of codes, such as the GPT-style model in VQGAN (Esser et al., 2021) or the BERT-style model in MaskGIT (Chang et al., 2022).
Besides differentiability, there are two interesting properties of the proposed SVQ scheme. First, according to our empirical study, we find that with large enough codebook size, we can freeze the randomly initialized codebook without sacrificing any performance, making SVQ potentially more efficient for training. Secondly, unlike lookup-based methods, SVQ allows for both positive and negative weights, enhancing the ability to add or subtract patterns to improve representational power.
Finally, despite its simplification, our experiments across multiple real-world benchmark datasets—including weather forecasting, traffic flow prediction, and video forecasting, demonstrate the superior performance of our SVQ method. In WeatherBench-S temperature forecasting benchmark, it improves the prior SOTA method by 7.9%. For video prediction across renowned datasets—Human3.6M, KTH, and KittiCaltech—SVQ not only achieves a consistent 9.4% reduction in MAE but also enhances perceptual quality, evidenced by a 17.3% lower LPIPS score. These results highlight SVQ’s versatility and superior accuracy in diverse spatial-temporal prediction scenarios.
2 Related Work
2.1 Recurrent-based forecasting model
The majority of spatio-temporal forecasting models leverage techniques such as Conv2D (Xu et al., 2018), Conv3D (Wang et al., 2019a), and attention (Liu et al., 2022b) for spatial modeling. Distinctions among these models primarily arise from how they incorporate temporal information. Recurrent-based models, exemplified by ConvLSTM (SHI et al., 2015), have been widely used to capture motion dynamics by iteratively processing multi-frame predictions. Variants like PredRNN (Wang et al., 2017) introduce the Spatio-Temporal LSTM (ST-LSTM) unit, integrating spatial appearances and temporal variations within a single memory pool. Further advancements include PredRNN++ (Wang et al., 2018) and PredRNNV2 (Wang et al., 2022b), which deepen the model and expand the receptive field through a cascading LSTM mechanism and a memory decoupling strategy, respectively.
2.2 Non-recurrent forecasting model
Despite the effectiveness of recurrent-based methods, they suffer from high computational cost caused by their inherent unparallelizable architecture. Recent efforts in spatio-temporal forecasting have shifted towards non-recurrent models that decouple the forecasting task from autoregressive processes. Notably, SimVPv1 (Gao et al., 2022) and SimVPv2 (Tan et al., 2022) separate spatial and temporal learning into distinct phases within an encoder-translator-decoder structure, consistently surpassing recurrent counterparts in video prediction tasks. Due to their similar architectures, we will collectively refer to them as the SimVP model in the remainder of our paper. TAU (Tan et al., 2023a) refines the architecture by incorporating a visual attention mechanism into the translator. OpenSTL (Tan et al., 2023b) further enhances the translator by MetaFormers.
2.3 Vector quantization and sparse coding
Instead of using continuous latent, VQVAE (van den Oord et al., 2017), a seminal work, integrates vector quantization (VQ) in learning discrete latent representations, typically assigning each vector to the nearest code in a codebook. Subsequent enhancements include Residual VQ (Zeghidour et al., 2022), which quantizes the residuals recursively, and Multi-headed VQ (Mama et al., 2021b), which adopts multiple heads for each vector. While effective, these methods often rely on a relatively small number of codes for representing continuous vectors. Addressing this, SCVAE (Xiao et al., 2023) employs sparse coding, allowing vectors to be represented through sparse linear combinations of multiple codes and achieving end-to-end training via the Learnable Iterative Shrinkage Thresholding Algorithm (LISTA) (Gregor & LeCun, 2010). However, a significant drawback of the sparse coding method using LISTA (dubbed SVQ-raw here) is its high computational complexity, which scales quadratically with codebook size.
Our work is closely related to a simultaneous research work (Tschannen et al., 2023), which employs an infinite cookbook with a linear layer for continuous vector quantization applied in image generation. Despite the similarity, our approach is largely inspired by sparse regression, which is clearly evidenced by our analysis. In particular, we provide a better theoretical foundation and address the challenges arising from spatio-temporal forecasting well.
3 Sparse Vector Quantization (SVQ)
In this section, we will first describe the mathematical foundation for sparse vector quantization, followed by the description of detailed implementation.
3.1 Vector quantization by sparse regression
Let be the set of codes. A typical vector quantization method assigns a data point to the nearest code in . The main problem with such an approach is that a significant part of the information in will be lost due to quantization. Sparse regression turns the code assignment problem into an optimization problem
where is the weight for combining codes to approximate . By introducing regularizer in the optimization problem, we effectively enforce to be associated with a small number of codes. Compared to classic VQ methods where codes have to be learned through clustering, according to (Chiu et al., 2022), it is sufficient to use randomly sampled vectors as codes as long as its number is large enough, thus avoiding the need of computing and adjusting codes. The theoretical guarantee of sparse regression is closely related to the property of subspace clustering, as revealed in Theorem 4.1.
As shown in Figure 2, the obvious downside of sparse regression for VQ is its high computational cost, as it needs to solve the optimization problem in (1) for EVERY data point. Below, we will show that sparse regression can be approximated by a two-layer MLP and a randomly fixed or learnable matrix, making it computationally attractive.
To solve the optimization problem (1), we consider the composite optimization method whose iteration is given as follows
where , is the th element of vector , and outputs if and otherwise. We consider the first step of the iteration where and have and the resulting output for is given as
By generalizing into another matrix , we have output vector exactly expressed as a matrix and a two-layer MLP over . We finally note that although it is convenient to form the codebook by randomly sampling vectors, we found empirically that tuning codebook does bring slight additional gains in some cases.
3.2 Detailed implementation
Architecture of backbone model. As depicted in Figure 3, we adopt SimVP as our backbone model, which encompasses an encoder for spatial feature extraction, a translator for temporal dependency learning, and a decoder for frame reconstruction. We integrate the quantization module between the encoder and translator. The input data is a 4D tensor , representing height (), width (), time steps (), and channels (). The encoder condenses into downsampled latent representation , maintaining temporal dimensionality while altering spatial and channel dimensions. This latent space, composed of tokens, each a -dimensional vector, undergoes vector quantization.
Quantization module. SVQ comprises a two-layer MLP and a huge codebook. The codebook is a randomly initialized matrix , where denotes the size of codebook. To achieve automatic selection of codes, we generate a weight matrix by nonlinear projection from latent representation . This projection is formally expressed as , wherein the comprises two linear layers and an intermediate ReLU activation function. Subsequently, the quantized output is obtained by computing the dot product of weight matrix and codebook matrix , a process that can be conceptualized as a selection operation as shown in Figure 3. To encourage sparsity within the generated weight matrix, we apply a Mean Absolute Error (MAE) loss to the output as a surrogate form of regularization.
Despite the potential feasibility of placing the quantization module either before or after the Translator, we note that post-Translator placement induces significant training instability in classic VQ, attributable to the non-differentiable nature of the straight-through estimator. Conversely, our SVQ module demonstrates consistent stability during training. For consistency, the quantization module is placed before the Translator in our primary experiments, with a comparative assessment of its alternative placement provided in Section B.1.
4 Efficient Utilize of Cookbook Using Sparse Regression
To understand the difference between sparse regression-based quantization scheme and clustering-based quantization scheme, we measure the number of codes that we need in order to approximate any vector within an unit ball with error less than . We denote this number by . As indicated by the theorem below, using spare regression, can be reduced from to , with for high dimensional vectors.
Theorem 4.1.
for the clustering-based method is at least , while for sparse regression can be , with
as long as the non-zero elements used by sparse regression is at least
Proof.
To bound for the clustering method, we use the covering number for a unit ball , requiring at least code vectors to approximate any vector within error . Considering , with , , and an -sparse unit vector, we have:
implying
with probability at least . Thus, . With the -sparse unit vector covering number bounded by , we obtain:
By choosing , we have
yielding , given .
As , it is easy to verify that , leading to . ∎
5 Experiments
We extensively evaluate SVQ on a wide range of real-world spatio-temporal forecasting datasets under the unified framework of OpenSTL (Tan et al., 2023b). Note that since SimVP holds leading performance across almost all benchmarks, it is adopted as our primary baseline. We begin by comparing the SimVP+SVQ model with state-of-the-art (SOTA) counterparts. Subsequently, we assess the capability of SVQ to enhance various backbones. Finally, we investigate the delicate balance between detail preservation and noise reduction, and conduct experiments with artificial noise injection to show the effect of SVQ on mitigating noise.
Moreover, we provide a comprehensive comparison between SVQ and existing VQ methods in Appendix B.3, evaluating their downstream prediction performance, codebook utilization, and computational efficiency.
Dataset. To evaluate the proposed SVQ, we conduct extensive experiments on five real-world spatio-temporal forecasting tasks, including weather (WeatherBench), traffic flow (TaxiBJ), human pose dynamics (Human3.6M), driving scenes (KittiCaltech), and human actions (KTH Action). More details about datasets are described in Appendix A.1.
Experimental details. In deployment, we found SVQ to be quite robust to codebook size, as its performance remains consistently strong when using a sufficiently large codebook. Therefore, we fix the codebook size at 10,000 for WeatherBench, TaxiBJ, and Human3.6M datasets, and at 6,000 for KittiCaltech and KTH datasets. The hidden dimension of nonlinear projection layer is fixed at 128. Our experiments are conducted on either 1 or 4 NVIDIA V100 32GB GPUs, with a total batch size of 16. More details about backbone architectures, VQ parameters, computational costs, and metrics are described in Appendix A.2, A.3, B.2, and A.5 respectively.
5.1 Benchmarks on various tasks
| Dataset | Variable | Temperature | Humidity | Wind Component | Total Cloud Cover | ||||
| Model | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | |
| WeatherBench-S | ConvLSTM(SHI et al., 2015) | 1.521 | 0.7949 | 35.146 | 4.012 | 1.8976 | 0.9215 | 0.0494 | 0.1542 |
| E3D-LSTM(Wang et al., 2019a) | 1.592 | 0.8059 | 36.534 | 4.100 | 2.4111 | 1.0342 | 0.0573 | 0.1529 | |
| PredRNN(Wang et al., 2017) | 1.331 | 0.7246 | 37.611 | 4.096 | 1.8810 | 0.9068 | 0.0550 | 0.1588 | |
| MIM(Wang et al., 2019b) | 1.784 | 0.8716 | 36.534 | 4.100 | 3.1399 | 1.1837 | 0.0573 | 0.1529 | |
| MAU(Chang et al., 2021) | 1.251 | 0.7036 | 34.529 | 4.004 | 1.9001 | 0.9194 | 0.0496 | 0.1516 | |
| PredRNN++(Wang et al., 2018) | 1.634 | 0.7883 | 35.146 | 4.012 | 1.8727 | 0.9019 | 0.0547 | 0.1543 | |
| PredRNN.V2(Wang et al., 2022b) | 1.545 | 0.7986 | 36.508 | 4.087 | 2.0072 | 0.9413 | 0.0505 | 0.1587 | |
| TAU(Tan et al., 2023a) | 1.162 | 0.6707 | 31.831 | 3.818 | 1.5925 | 0.8426 | 0.0472 | 0.1460 | |
| SimVP (w/o VQ)(Tan et al., 2022) | 1.105 | 0.6567 | 31.332 | 3.776 | 1.4996 | 0.8145 | 0.0466 | 0.1469 | |
| SimVP+SVQ (Frozen) | 1.023 | 0.6131 | 30.863 | 3.661 | 1.4337 | 0.7861 | 0.0456 | 0.1456 | |
| SimVP+SVQ (Learnable) | 1.018 | 0.6109 | 30.611 | 3.657 | 1.4186 | 0.7858 | 0.0458 | 0.1463 | |
| Improvement | 7.9% | 7.0% | 2.3% | 3.2% | 5.4% | 3.5% | 2.1% | 0.9% | |
| WeatherBench-M | Variable | Temperature | Humidity | Wind U Component | Wind V Component | ||||
| ConvLSTM(SHI et al., 2015) | 6.303 | 1.7695 | 368.15 | 13.490 | 30.002 | 3.8923 | 30.789 | 3.8238 | |
| PredRNN(Wang et al., 2017) | 5.596 | 1.6411 | 354.57 | 13.169 | 27.484 | 3.6776 | 28.973 | 3.6617 | |
| MIM(Wang et al., 2019b) | 7.515 | 1.9650 | 408.24 | 14.658 | 35.586 | 4.2842 | 36.464 | 4.2066 | |
| MAU(Chang et al., 2021) | 5.628 | 1.6810 | 363.36 | 13.503 | 27.582 | 3.7409 | 27.929 | 3.6700 | |
| PredRNN++(Wang et al., 2018) | 5.647 | 1.6433 | 363.15 | 13.246 | 28.396 | 3.7322 | 29.872 | 3.7067 | |
| PredRNN.V2(Wang et al., 2022b) | 6.307 | 1.7770 | 368.52 | 13.594 | 29.833 | 3.8870 | 31.119 | 3.8406 | |
| TAU(Tan et al., 2023a) | 4.904 | 1.5341 | 342.63 | 12.801 | 24.719 | 3.5060 | 25.456 | 3.4723 | |
| SimVP (w/o VQ)(Tan et al., 2022) | 4.833 | 1.5246 | 340.06 | 12.738 | 24.535 | 3.4882 | 25.232 | 3.4509 | |
| SimVP+SVQ (Frozen) | 4.427 | 1.4160 | 360.15 | 12.445 | 23.915 | 3.4078 | 24.968 | 3.4117 | |
| SimVP+SVQ (Learnable) | 4.433 | 1.4164 | 360.53 | 12.449 | 23.908 | 3.4060 | 24.983 | 3.4095 | |
| Improvement | 8.4% | 7.1% | 5.9% | 2.3% | 2.6% | 2.4% | 1.0% | 1.2% | |
| Dataset | Human3.6M | KittiCaltech | KTH | |||||||||
| Metric | MAE | SSIM | PSNR | LPIPS | MAE | SSIM | PSNR | LPIPS | MAE | SSIM | PSNR | LPIPS |
| ConvLSTM(SHI et al., 2015) | 1583.3 | 0.9813 | 33.40 | 0.03557 | 1583.3 | 0.9345 | 27.46 | 0.08575 | 445.5 | 0.8977 | 26.99 | 0.26686 |
| E3D-LSTM(Wang et al., 2019a) | 1442.5 | 0.9803 | 32.52 | 0.04133 | 1946.2 | 0.9047 | 25.45 | 0.12602 | 892.7 | 0.8153 | 21.78 | 0.48358 |
| PredNet(Lotter et al., 2017) | 1625.3 | 0.9786 | 31.76 | 0.03264 | 1568.9 | 0.9286 | 27.21 | 0.11289 | 783.1 | 0.8094 | 22.45 | 0.32159 |
| PhyDNet(Guen & Thome, 2020) | 1614.7 | 0.9804 | 39.84 | 0.03709 | 2754.8 | 0.8615 | 23.26 | 0.32194 | 765.6 | 0.8322 | 23.41 | 0.50155 |
| MAU(Chang et al., 2021) | 1577.0 | 0.9812 | 33.33 | 0.03561 | 1800.4 | 0.9176 | 26.14 | 0.09673 | 471.2 | 0.8945 | 26.73 | 0.25442 |
| MIM(Wang et al., 2019b) | 1467.1 | 0.9829 | 33.97 | 0.03338 | 1464.0 | 0.9409 | 28.10 | 0.06353 | 380.8 | 0.9025 | 27.78 | 0.18808 |
| PredRNN(Wang et al., 2017) | 1458.3 | 0.9831 | 33.94 | 0.03245 | 1525.5 | 0.9374 | 27.81 | 0.07395 | 380.6 | 0.9097 | 27.95 | 0.21892 |
| PredRNN++(Wang et al., 2018) | 1452.2 | 0.9832 | 34.02 | 0.03196 | 1453.2 | 0.9433 | 28.02 | 0.13210 | 370.4 | 0.9124 | 28.13 | 0.19871 |
| PredRNN.V2(Wang et al., 2022b) | 1484.7 | 0.9827 | 33.84 | 0.03334 | 1610.5 | 0.9330 | 27.12 | 0.08920 | 368.8 | 0.9099 | 28.01 | 0.21478 |
| TAU(Tan et al., 2023a) | 1390.7 | 0.9839 | 34.03 | 0.02783 | 1507.8 | 0.9456 | 27.83 | 0.05494 | 421.7 | 0.9086 | 27.10 | 0.22856 |
| DMVFN (Hu et al., 2023) | - | - | - | - | 1531.1 | 0.9314 | 26.95 | 0.04942 | 413.2 | 0.8976 | 26.65 | 0.12842 |
| SimVP (w/o VQ)(Tan et al., 2022) | 1441.0 | 0.9834 | 34.08 | 0.03224 | 1507.7 | 0.9453 | 27.89 | 0.05740 | 397.1 | 0.9065 | 27.46 | 0.26496 |
| SimVP+SVQ (Frozen) | 1264.9 | 0.9851 | 34.07 | 0.02380 | 1408.6 | 0.9469 | 28.10 | 0.05535 | 364.6 | 0.9109 | 27.28 | 0.20988 |
| SimVP+SVQ (Learnable) | 1265.1 | 0.9851 | 34.06 | 0.02367 | 1414.9 | 0.9458 | 28.10 | 0.05776 | 360.2 | 0.9116 | 27.37 | 0.20658 |
| Improvement | 12.2% | 0.2% | 0.0% | 26.2% | 6.6% | 0.2% | 0.8% | 3.6% | 9.3% | 0.6% | 0.3% | 22.0% |
We explore both fixed (frozen) and learnable versions of SVQ on various forecasting tasks. Interestingly, we find that when codebook size is large, the performance of frozen, randomly-initialized codebook is on par with carefully learned codebook. Our comparison baselines consist of two categories: 1) Non-recurrent models including SimVP (Tan et al., 2022) and TAU (Tan et al., 2023a); 2) Recurrent-based models including ConvLSTM (SHI et al., 2015), PredNet (Lotter et al., 2017), PredRNN (Wang et al., 2017), PredRNN++ (Wang et al., 2018), MIM (Wang et al., 2019b), E3D-LSTM (Wang et al., 2019a), PhyDNet (Guen & Thome, 2020), MAU (Chang et al., 2021), PredRNNv2 (Wang et al., 2022b), and DMVFN (Hu et al., 2023). Baseline results are copied from the original OpenSTL paper (Tan et al., 2023b). To preclude ambiguity, we select the best MetaFormer of SimVP for each dataset, detailed in Appendix A.2.
We present benchmark results of WeatherBench and three video prediction datasets (Human3.6M, KTH, and KittiCaltech) in Tables 1 and 2, respectively. Due to page limit, the benchmark results of TaxiBJ dataset are provided in Appendix B.4. These datasets have different characteristics. WeatherBench and TaxiBJ are macro forecasting tasks with low-frequency collection (30min or 1-6h). Human3.6M features subtle differences between frames as low-frequency signals. KittiCaltech poses challenges due to its rapidly changing backgrounds and limited training data. The KTH dataset, on the other hand, tests the capacity for long-horizon forecasting, requiring the prediction of 20 future frames based on 10 observed frames.
However, despite the distinct characteristics among datasets, a common thread is the need for improved noise reduction coupled with enhanced representational capabilities, which can universally benefit their respective forecasting tasks. Notably, the SimVP+SVQ model achieves either the best or comparable performance across all datasets. For instance, on the WeatherBench-S temperature dataset, SVQ significantly improves the best baseline by 7.9% (1.105 → 1.018). On these three popular video prediction tasks, SVQ not only delivers a reduction in forecasting errors (average 9.4% decrease in MAE), but also significantly improves subjective image quality (average 17.3% decrease in LPIPS). Additional visualizations of forecasting samples can be found in Appendix D.
5.2 Boosting performance as a versatile plug-in
In this section, SVQ serves as a versatile plug-in module applicable to various MetaFormers (Yu et al., 2022a). The adopted MetaFormers can be categorized into three types. 1) CNN-based: SimVPv1(IncepU) (Gao et al., 2022), SimVPv2(gSTA) (Tan et al., 2022), ConvMixer (Trockman & Kolter, 2023), ConvNeXt (Liu et al., 2022a), HorNet (Rao et al., 2022), and MogaNet (Li et al., 2022). 2) Transformer-based: ViT (Dosovitskiy et al., 2021), Swin Transformer (Liu et al., 2021), Uniformer (Li et al., 2023), Poolformer (Yu et al., 2022b), and VAN (Guo et al., 2023). 3) MLP-based: MLPMixer (Tolstikhin et al., 2021). We conduct experiments on WeatherBench-S temperature dataset because it is lightweight and fast for training.
As shown in Table 3, SVQ consistently improves the performance across all MetaFormers, showcasing its universality across diverse backbone types. We observe an average reduction in MSE and MAE by 4.8% and 6.0%, respectively. In detail, SVQ leads to an average MSE reduction of 4.1% for CNN-based backbones, 5.1% for transformer-based, and 10.7% for MLP-based. The more pronounced enhancement in transformer-based and MLP-based models indicates that our approach is especially effective with architectures that prioritize global interactions. Notably, SimVPv2(gSTA) is the best backbone, while our SVQ further improves it by 7.9%. These findings also aligns with our motivation that mitigating noise in the learning process significantly benefits spatio-temporal forecasting, irrespective of model architecture. By integrating SVQ to constrain the diversity of predicted patterns and cut out noise, researchers can focus on crafting high-quality and general base models.
| MetaFormer | MSE | MAE | ||
| w/o SVQ | w SVQ | w/o SVQ | w SVQ | |
| SimVPv1(IncepU)(Gao et al., 2022) | 1.238 | 1.216 | 0.7037 | 0.6831 |
| SimVPv2(gSTA)(Tan et al., 2022) | 1.105 | 1.018 | 0.6567 | 0.6109 |
| ConvMixer(Trockman & Kolter, 2023) | 1.267 | 1.257 | 0.7073 | 0.6780 |
| ConvNeXt(Liu et al., 2022a) | 1.277 | 1.159 | 0.7220 | 0.6568 |
| HorNet(Rao et al., 2022) | 1.201 | 1.130 | 0.6906 | 0.6472 |
| MogaNet(Li et al., 2022) | 1.152 | 1.067 | 0.6665 | 0.6271 |
| ViT(Dosovitskiy et al., 2021) | 1.146 | 1.111 | 0.6712 | 0.6375 |
| Swin(Liu et al., 2021) | 1.143 | 1.088 | 0.6735 | 0.6320 |
| Uniformer(Li et al., 2023) | 1.204 | 1.110 | 0.6885 | 0.6400 |
| Poolformer(Yu et al., 2022b) | 1.156 | 1.097 | 0.6715 | 0.6297 |
| VAN(Guo et al., 2023) | 1.150 | 1.083 | 0.6803 | 0.6342 |
| MLP-Mixer(Tolstikhin et al., 2021) | 1.255 | 1.120 | 0.7011 | 0.6455 |
| Average improvement | 4.8% | 6.0% | ||
5.3 Delicate balance between detail preservation and noise reduction
Investigating vector quantization’s (VQ) role in spatio-temporal forecasting, we evaluated several cutting-edge VQ methods akin to the SVQ framework, implemented as plug-in modules coordinated with the primary model. Table 4 shows that SVQ markedly improves forecasting as a plug-in, whereas other VQ approaches generally lead to heightened prediction errors. Enhanced detail retention within the representational capacity is linked to lower forecasting errors. Classic VQ methods suffer from notable information losses, as evidenced by a higher MSE of 1.854. In contrast, both residual VQ and grouped residual VQ outperform traditional VQ with lower MSEs of 1.213 and 1.174, respectively, affirming their ability to preserve intricate details.
It is commonly understood that the size of the codebook in clustering-based VQ is critical: larger codebooks capture more details, whereas smaller ones enhance noise reduction. To explore this trade-off, we analyzed how the codebook size influences the prediction MSE in grouped residual VQ. As Figure 4 indicates, MSE initially decreases but increases with overly large codebooks, echoing findings from (Yu et al., 2023b) that an excessively large codebook may degrade image generation performance. This underscores the necessity for dataset-specific tuning in clustering-based VQ approaches. In contrast, SVQ naturally balances detail preservation and noise reduction using sparse regression, negating the need for such fine-tuning.The codebook size in our approach exhibits a low-maintenance profile: performance remains solid across various codebook sizes, suggesting that utilizing a default large codebook can yield robust results without the need for extensive tuning. We are not suggesting that our SVQ outperforms other VQ methods in general image generation tasks, as this is beyond the scope of our current objective. Rather, we emphasize SVQ’s effectiveness as a noise reduction tool that directly enhances real-world spatio-temporal forecasting tasks, while the application to general image generation remains a topic for future exploration.
| Method | MSE | MAE |
| Baseline (SimVP w/o VQ) | 1.105 | 0.6567 |
| VQ(van den Oord et al., 2017) | 1.854 | 0.8963 |
| Residual VQ(Zeghidour et al., 2022) | 1.213 | 0.6910 |
| Grouped Residual VQ(Yang et al., 2023) | 1.174 | 0.6747 |
| Multi-headed VQ(Mama et al., 2021b) | 1.211 | 0.6994 |
| Residual VQ (Stochastic)(Lee et al., 2022) | 1.888 | 0.9237 |
| Residual Finite Scalar Quantization(Mama et al., 2021a) | 1.319 | 0.7505 |
| Lookup Free Quantization (LFQ)(Yu et al., 2023a) | 2.988 | 1.1103 |
| Residual LFQ(Yu et al., 2023a) | 1.281 | 0.7281 |
| SVQ-raw(Xiao et al., 2023) | 1.123 | 0.6456 |
| SVQ | 1.018 | 0.6109 |
5.4 Robustness to artificial noise injection
To clearly demonstrate the noise reduction effect of SVQ, we conduct an experiment by adding artificial noise to the training data. Table 5 shows that the SVQ-equiped model experiences a much lower rise in MSE and MAE relative to the model without SVQ, regardless of the proportion of injected noise. For instance, when the proportion of injected noise is 10%, the MSE of the model without SVQ rises by 25.4%, whereas the model with SVQ shows a modest increase of just 3.2%. These results explicitly confirm that our method effectively mitigates noise by constraining latent patterns via quantization.
| Noise proportion | MSE | MAE | ||
| w/o SVQ | w SVQ | w/o SVQ | w SVQ | |
| =0 | 1.105 | 1.018 | 0.6567 | 0.6109 |
| =10% | 1.386(+25.4%) | 1.051(+3.2%) | 0.7702(+17.3%) | 0.6269(+2.6%) |
| =20% | 1.554(+40.7%) | 1.196(+17.5%) | 0.8282(+26.1%) | 0.6710(+9.8%) |
| =30% | 1.750(+58.4%) | 1.255(+23.2%) | 0.8821(+34.3%) | 0.6953(+13.8%) |
| =40% | 2.081(+88.3%) | 1.568(+54.0%) | 1.0031(+52.7%) | 0.7973(+30.5%) |
| =50% | 2.646(+139.4%) | 1.529(+50.2%) | 1.1339(+72.7%) | 0.7881(+29.0%) |
5.5 Ablation study
We conduct a series of ablation studies on WeatherBench-S temperature dataset to understand the contribution of important designs based on the default setting: SVQ with 10000 codebook size, learnable codebook, and MAE loss.
Codebook size and learnability. We compare the effect of codebook size when codebook is learnable or frozen (Table 6). The results indicate that that increasing codebook size gradually improves performance. When significantly increasing codebook size to a large value, such as 10000, the performance difference between frozen and learnable codebook narrows, evidenced by a mere 0.5% (1.023 → 1.018). We observe that enlarging the codebook ensures comprehensive latent space coverage through random codes, negating the need for meticulous learning and allowing random initializations to rival learned counterparts. Additionally, our tailored SVQ structure surpasses alternative designs, including single-layer, bucket-shaped, and post-ReLU variants, while maintaining a similar parameter count.
| Learnability | Projection dim | Codebook size | MSE | MAE |
| Frozen | 128 | 10 | 1.070 | 0.6227 |
| 128 | 1000 | 1.044 | 0.6198 | |
| 128 | 10000 | 1.023 | 0.6131 | |
| Learnable | 128 | 10 | 1.060 | 0.6194 |
| 128 | 1000 | 1.048 | 0.6182 | |
| 128 | 10000 | 1.018 | 0.6109 | |
| 1280(Bucket-shape) | 1280 | 1.035 | 0.6149 | |
| None(One-layer) | 10000 | 1.043 | 0.6144 | |
| 128(Post-ReLU) | 10000 | 1.032 | 0.6136 |
Self-learned sparse regression structure. The original SimVP model adopts MSE as prediction loss. We individually replace it with MAE loss and add the SVQ module. The results show that the joint use of SVQ and MAE loss is crucial for significantly improving the model’s performance (Table 7). We suggest that the sparsity of the weight matrix impacts vector representation learning and use kurtosis to quantify this after normalizing . Figure 5 demonstrates that both a learnable codebook and MAE loss contribute to increased sparsity. Analyzing four codebook initialization methods, in both learnable and fixed settings (Table 8), we find that a learnable codebook promotes sparsity irrespective of the initialization, indicating that sparsity is a self-learned property that enhances intermediate representation learning.
| Module | MSE | MAE |
| SimVP (MSE loss) | 1.105 | 0.6567 |
| SimVP (MAE loss) | 1.126 | 0.6509 |
| SimVP+SVQ (Learnable, MSE loss) | 1.099 | 0.6527 |
| SimVP+SVQ (Learnable, MAE loss) | 1.018 | 0.6109 |
| Initialization | Learnability | MSE | MAE | Kurtosis |
| kaiming uniform | Frozen | 1.023 | 0.6131 | 1.596 |
| Learnable | 1.018 | 0.6109 | 7.213 | |
| sparse(sparsity=0.9) | Frozen | 1.050 | 0.6183 | 4.165 |
| Learnable | 1.034 | 0.6160 | 41.558 | |
| trunc normal | Frozen | 1.049 | 0.6166 | 1.582 |
| Learnable | 1.031 | 0.6161 | 4.236 | |
| orthogonal | Frozen | 1.034 | 0.6170 | 1.561 |
| Learnable | 1.030 | 0.6131 | 35.774 |
6 Conclusions
In this work, we present Sparse regression-based Vector Quantization (SVQ), a concise yet effective method for spatiotemporal forecasting enhancement. SVQ elegantly balances detail preservation and noise reduction using sparse regression, realized via a two-layer MLP and a codebook, offering computational efficiency and full differentiability. Tested across diverse benchmarks, from weather to traffic and video prediction, SVQ consistently outperforms existing models, setting new performance standards without complex priors. Its differentiability and seamless integration with baseline models highlight SVQ as a significant advancement for efficient and effective spatio-temporal forecasting.
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Supplemental Materials
The supplementary material for our work A Differentiable Sparse Vector Quantization (SVQ) for Spatio-Temporal Forecasting is organized as follows: Section A provides implementation details of SimVP model and VQ methods. Section B present extended quantitative results. Section C examines the effect of SVQ on latent representation. Finally, Section D shows additional qualitative results of forecasting samples and errors.
Appendix A Implementation details
A.1 Dataset details
The details of the benchmark datasets are summarized as follows: WeatherBench (Rasp et al., 2020) and TaxiBJ (Zhang et al., 2017) are two macro forecasting tasks. These two datasets are collected at low frequencies (30min or 1-6h) and exhibit stable states, which is especially challenging for the model to capture subtle changes. Human3.6M (Ionescu et al., 2014), KittiCaltech (Geiger et al., 2013; Dollár et al., 2009), and KTH Action (Schüldt et al., 2004) are three popular video prediction tasks. Due to the diversity and complexity of real-world videos, these datasets have totally different characteristics. A summary of dataset statistics is provided in Table 9.
| Dataset | Size | Seq. Len. | Img. Shape | Interval | ||
| train | test | in | out | H W C | ||
| WeatherBench-S | 2,167 | 706 | 12 | 12 | 32 64 1 | 1 hour |
| WeatherBench-M | 54,019 | 2,883 | 4 | 4 | 32 64 4 | 6 hour |
| TaxiBJ | 20,461 | 500 | 4 | 4 | 32 32 2 | 30 min |
| KittiCaltech | 3,160 | 3,095 | 10 | 1 | 128 160 3 | Frame |
| Human3.6M | 73,404 | 8,582 | 4 | 4 | 256 256 3 | Frame |
| KTH Action | 4,940 | 3,030 | 10 | 20 | 128 128 1 | Frame |
A.2 Architecture configuration of SimVP
Table 10 reports the architectures of SimVP on all datasets. We select the best MetaFormer to replace the translator module based on OpenSTL benchmarks111https://openstl.readthedocs.io/en/latest/model_zoos/video_benchmarks.html222https://openstl.readthedocs.io/en/latest/model_zoos/weather_benchmarks.html333https://openstl.readthedocs.io/en/latest/model_zoos/traffic_benchmarks.html. The parameters remain unchanged, following the original configurations. It should be noted that, due to the reproducibility issue of ConvNeXt on the TaxiBJ dataset, we have opted to utilize gSTA as our backbone model.
| Dataset | MetaFormer (Translator) | spatio_kernel | hid_S | hid_T | N_T | N_S | drop_path | LR scheduler |
| WeatherBench-S temperature | gSTA | enc=3, dec=3 | 32 | 256 | 8 | 2 | 0.1 | cosine |
| WeatherBench-S humidity | Swin | enc=3, dec=3 | 32 | 256 | 8 | 2 | 0.2 | cosine |
| WeatherBench-S wind component | Swin | enc=3, dec=3 | 32 | 256 | 8 | 2 | 0.2 | cosine |
| WeatherBench-S total cloud cover | gSTA | enc=3, dec=3 | 32 | 256 | 8 | 2 | 0.1 | cosine |
| WeatherBench-M | MogaNet | enc=3, dec=3 | 32 | 256 | 8 | 2 | 0.1 | cosine |
| TaxiBJ | gSTA | enc=3, dec=3 | 32 | 256 | 8 | 2 | 0.1 | cosine |
| Human3.6M | gSTA | enc=3, dec=3 | 64 | 512 | 6 | 4 | 0.1 | cosine |
| KTH | IncepU | enc=3, dec=3 | 64 | 256 | 6 | 2 | 0.1 | onecycle |
| KittiCaltech | gSTA | enc=3, dec=3 | 64 | 256 | 6 | 2 | 0.2 | onecycle |
A.3 Parameters of compared VQ methods
In Table 4, we compared SVQ with a series of well-known VQ methods. These methods are reproduced using the source code provided on the GitHub repository444https://github.com/lucidrains/vector-quantize-pytorch/tree/master. We maintain the parameters consistent with the recommended settings to guarantee performance, as detailed in Table 11. Importantly, to maintain comparable complexity among the methods, we enable a shared codebook for all quantizers in Residual VQ, as well as across the heads in Multi-headed VQ. Moreover, we also conduct an extensive additional experiment for VQ methods using the same codebook size in Section B.3.
| Vector quantization method | codebook_size | num_quantizers | groups | heads | shared_codebook | Specific parameters |
| VQ | 512 | - | - | - | - | - |
| Residual VQ | 1024 | 8 | - | - | - | |
| Grouped Residual VQ | 1024 | 8 | 2 | - | - | |
| Multi-headed VQ | 1024 | - | - | 8 | - | |
| Residual VQ (Stochastic) | 1024 | 8 | - | - | stochastic_sample_codes=True | |
| Residual Finite Scalar Quantization | - | 8 | - | - | - | levels=[8, 5, 5, 3] |
| Lookup Free Quantization (LFQ) | 8192 | - | - | - | - | entropy_loss_weight=0.1 |
| Residual LFQ | 256 | 8 | - | - | - | - |
A.4 Evaluation of perplexity
In contrast to other VQ methods that rely on a single code, our SVQ generates multiple regression weights to merge several codes. To evaluate its perplexity, we first normalize the regression weights and then convert them into binary form using a threshold set at times the standard deviation, where serves as the threshold value. We utilize two thresholds (2 and 3) to obtain reasonable perplexity.
A.5 Metrics
Forecasting accuracy is evaluated using mean squared error (MSE) and mean absolute error (MAE), while the image quality of predicted frames is assessed using structural similarity index measure (SSIM) (Wang et al., 2004), peak signal-to-noise ratio (PSNR), and learned perceptual image patch similarity (LPIPS) (Zhang et al., 2018). The training process is early stopped with a patience of 10, and the models with the minimal loss are saved for subsequent evaluation.
Appendix B Additional quantitative results
B.1 Position of quantization module
To illustrate how the position of quantization module affects representation learning, we consider two designs as shown in Figure 6. In our main experiments, we adopt the first design where quantization is performed before the translator. In this section, we also investigate an alternative design where quantization is performed after the translator. Interestingly, we found that the classic VQ method (van den Oord et al., 2017) suffers pronounced instability and codebook collapse issues under the latter design, as shown in Figure 7. It is essential to highlight that the backbones without VQ maintain their MSE within the acceptable range of approximately 1 to 2 (refer to Table 3). Yet, integrating classic VQ causes a substantial rise in MSE values, exceeding 10 for different backbones—a level considered excessively high. We hypothesize that this instability is attributed to the non-differentiability of the straight-through estimator, which introduces errors into the gradient flow for preceding modules. In contrast, our SVQ module never encounters such issues and remains highly stable throughout training. To maintain the integrity of all VQ methods, we opt for the first design in our principal manuscript, wherein quantization is executed preceding the translator module.
B.2 Computational cost
Table 12 presents the computational costs of SVQ module and forecasting models. It shows that recurrent-based models have significantly higher FLOPS requirements, while non-recurrent models are more efficient. The proposed SVQ module is not only effective but also computationally cheap. Across all datasets, SVQ only slightly adds the number of parameters and FLOPS. The computational burden of SimVP+SVQ remains significantly smaller than recurrent-based models.
| Model type | Dataset | Human3.6M | KTH | KittiCaltech | WeatherBench-S | TaxiBJ | |||||
| Model | Params | FLOPS | Params | FLOPS | Params | FLOPS | Params | FLOPS | Params | FLOPS | |
| Recurrent-based | ConvLSTM | 15.5M | 347.0G | 14.9M | 1368.0G | 15.0M | 595.0G | 14.98M | 136G | 14.98M | 20.74G |
| E3D-LSTM | 60.9M | 542.0G | 53.5M | 217.0G | 54.9M | 1004G | 51.09M | 169G | 50.99M | 98.19G | |
| PredNet | 12.5M | 13.7G | 12.5M | 3.4G | 12.5M | 12.5M | - | - | 12.5M | 0.85G | |
| PhyDNet | 4.2M | 19.1G | 3.1M | 93.6G | 3.1M | 40.4G | 3.09M | 36.8G | 3.09M | 5.60G | |
| MAU | 20.2M | 105.0G | 20.1M | 399.0G | 24.3M | 172.0G | 5.46M | 39.6G | 4.41M | 6.02G | |
| MIM | 47.6M | 1051.0G | 39.8M | 1099.0G | 49.2M | 1858G | 37.75M | 109G | 37.86M | 64.10G | |
| PredRNN | 24.6M | 704.0G | 23.6M | 2800.0G | 23.7M | 1216G | 23.57M | 278G | 23.66M | 42.40G | |
| PredRNN++ | 39.3M | 1033.0G | 38.3M | 4162.0G | 38.5M | 1803G | 38.31M | 413G | 38.40M | 62.95G | |
| PredRNN.V2 | 24.6M | 708.0G | 23.6M | 2815.0G | 23.8M | 1223G | 23.59M | 279G | 23.67M | 42.63G | |
| DMVFN | - | - | 3.5M | 0.88G | 3.6M | 1.2G | - | - | 3.54M | 0.057G | |
| Non-recurrent | TAU | 37.6M | 182.0G | 15.0M | 73.8G | 44.7M | 80.0G | 12.22M | 6.70G | 9.55M | 2.49G |
| SimVP (w/o VQ) | 28.8M | 146.0G | 12.2M | 62.8G | 15.6M | 96.3G | 12.76M | 7.01G | 7.84M | 2.08G | |
| SimVP+SVQ | 30.7M | 178.0G | 13.3M | 110.0G | 16.8M | 156G | 14.37M | 16.8G | 9.45M | 3.72G | |
B.3 Comprehensive comparison of VQ methods using the same codebook size
To make a fair comparison, we extend Table 4 by setting the codebook size to the same value (1024) for compared VQ methods. They are comprehensively evaluated from different aspects including downstream performance (prediction MSE), codebook usage (perplexity), and computational complexity (FLOPS, inference FPS, and training time per epoch). The quantitative results are shown in Table 13. The convergence performance is shown in Figure 8, where SVQ quickly converges to the lowest prediction error and satisfactory utilization of the codebook. Residual VQ with stochastic sampling has the highest codebook usage. However, its prediction MSE is worse than residual VQ without stochastic sampling. This demonstrates that codebook usage does not guarantee better downstream performance. SVQ generally outperforms the other VQ methods in computational efficiency, due to the simplified approximation described in Section 3.1.
| Vector quantization method | Prediction MSE | Perplexity | FLOPS | Inference FPS | Training time per epoch(min) |
| VQ | 1.8544 | 51.95 | 7.207G | 21.1 | 7.11 |
| Residual VQ | 1.2131 | 142.47 | 8.616G | 7.7 | 13.25 |
| Residual VQ (Stochastic) | 1.8882 | 817.91 | 8.616G | 8.1 | 17.27 |
| Grouped Residual VQ | 1.1737 | 132.57 | 8.616G | 4.8 | 19.98 |
| Multi-headed VQ | 1.2113 | 16.36 | 8.717G | 6.2 | 13.15 |
| SVQ (freeze) | 1.0393 | 335.72(=3)/438.39(=2) | 8.037G | 24.6 | 7.27 |
| SVQ (Learnable) | 1.0403 | 246.44(=3)/331.41(=2) | 8.037G | 24.9 | 7.30 |
B.4 Benchmark on TaxiBJ dataset
| Model | MSE | MAE | SSIM | PSNR |
| ConvLSTM(SHI et al., 2015) | 0.3358 | 15.32 | 0.9836 | 39.45 |
| E3D-LSTM(Wang et al., 2019a) | 0.3427 | 14.98 | 0.9842 | 39.64 |
| PhyDNet(Guen & Thome, 2020) | 0.3622 | 15.53 | 0.9828 | 39.46 |
| PredNet(Lotter et al., 2017) | 0.3516 | 15.91 | 0.9828 | 39.29 |
| PredRNN(Wang et al., 2017) | 0.3194 | 15.31 | 0.9838 | 39.51 |
| MIM(Wang et al., 2019b) | 0.3110 | 14.96 | 0.9847 | 39.65 |
| MAU(Chang et al., 2021) | 0.3268 | 15.26 | 0.9834 | 39.52 |
| DMVFN(Hu et al., 2023) | 3.3954 | 45.52 | 0.8321 | 31.14 |
| PredRNN++(Wang et al., 2018) | 0.3348 | 15.37 | 0.9834 | 39.47 |
| PredRNN.V2(Wang et al., 2022b) | 0.3834 | 15.55 | 0.9826 | 39.49 |
| TAU(Tan et al., 2023a) | 0.3108 | 14.93 | 0.9848 | 39.74 |
| SimVP (w/o VQ)(Tan et al., 2022) | 0.3246 | 15.03 | 0.9844 | 39.71 |
| SimVP+SVQ (Frozen) | 0.3171 | 14.68 | 0.9848 | 39.83 |
| SimVP+SVQ (Learnable) | 0.3191 | 14.64 | 0.9849 | 39.86 |
| Improvement | 1.7% | 2.6% | 0.1% | 0.4% |
B.5 Ablation of frozen module
The SVQ module consists of a two-layer MLP and a large codebook. The MLP can be seen as a projection from the input vector to the regression weights. In Section 5.5, we have already examined the impact of freezing the codebook on forecasting performance. To further investigate the effect of freezing the two-layer MLP, we conducted an ablation study in this section. The results, presented in Table 15, show that freezing the codebook only has a slight impact on forecasting performance, while freezing the MLP significantly impairs the performance.
| Frozen module | None (All learnable) | Codebook | Two-layer MLP projection | Both |
| MSE | 1.018 | 1.023 | 1.060 | 1.093 |
| MAE | 0.6109 | 0.6131 | 0.6194 | 0.6387 |
Appendix C Effect of SVQ on latent representation
We investigated the impact of SVQ on the sparsity of regression weights in Figure 5. To further investigate its effect on the latent representation, we compare the distribution of batch tensors before and after applying SVQ. We transform the tensors into normalized vectors and estimate their density distributions. As depicted in Figure 9, the representation after SVQ demonstrates a more compact distribution, indicating improved robustness to noise. The results further prove that SVQ can enhance forecasting performance by effectively handling noise in the data.
Figures 10, 11, 13, 14, and 12 present the comparison of latent feature maps before and after applying SVQ. These figures illustrate that the difference between foreground and background in the feature maps increases after SVQ. For example, in the KittiCaltech dataset, a clear distinction is observed between road conditions and sky (Figure 10). Similarly, in the WeatherBench-S temperature dataset, distinctive regions are identified between high and low latitudes (Figure 11). These findings suggest that SVQ helps in enhancing the discriminative power of the latent representations, which in turn contributes to improved downstream forecasting performance.
Appendix D Additional qualitative results
D.1 Forecasting errors on WeatherBench and TaxiBJ datasets
D.2 Additional forecasting samples